### v2

# Designing AI for Prosecutorial Governance: Case Prioritization and Statutory Oversight in Mexico

Abstract
Prosecutors across Mexico face growing backlogs due to high caseloads and limited institutional capacity. This paper presents a machine learning (ML) system co-developed with the Zacatecas State Prosecutor’s Office to support internal case triage. Focusing on the Módulo de Atención Temprana (MAT)—the unit responsible for intake and early-stage case resolution—we train classification models on administrative data from the state’s digital case management system (PIE) to predict which open cases are likely to finalize within six months. The model generates weekly ranked lists of 300 cases to assist prosecutors in identifying actionable files. Using historical data from 2014 to 2024, we evaluate model performance under real-time constraints, finding that Random Forest classifiers achieve a mean Precision@300 of 0.74. The system emphasizes interpretability and operational feasibility, and we will test it via a randomized controlled trial. Our results suggest that data-driven prioritization can serve as a low-overhead tool for improving prosecutorial efficiency without disrupting existing workflows.
Keywords artificial intelligence $\cdot$ digital government $\cdot$ criminal justice $\cdot$ algorithmic governance $\cdot$ case prioritization $\cdot$ public sector AI $\cdot$ decision support systems $\cdot$ Mexico
1 Introduction
Digital-era governance has fundamentally transformed how public institutions manage information, deliver services, and exercise administrative discretion (Dunleavy et al., 2006; Fountain, 2001). In criminal justice systems, where decisions carry profound consequences for individuals and communities, this transformation presents both opportunities and challenges. Algorithmic systems can help institutions process large caseloads more efficiently, but they also raise concerns about transparency, accountability, and the preservation of professional judgment (Janssen and Kuk, 2016; Veale et al., 2018). As governments increasingly deploy artificial intelligence (AI) in high-stakes domains, understanding how to design, implement, and oversee these systems becomes essential for both researchers and practitioners (Wirtz et al., 2019; Bullock et al., 2020).
This paper contributes to the growing literature on AI in government by presenting a case study of algorithmic decision support in a Mexican state prosecutor’s office. We describe the design, implementation, and governance of a machine learning system developed to address two interrelated institutional challenges: (1) supporting prosecutors in identifying cases likely to be resolved in the near term, and (2) providing systematic oversight of cases that may have exceeded statutory time limits. The system thus serves a dual function—enhancing operational efficiency while simultaneously enabling legal accountability. This dual-purpose design reflects a broader lesson for public sector AI: algorithmic tools need not be limited to efficiency gains but can also be structured to support institutional oversight and democratic governance.
Our analysis responds to calls for empirical research on AI implementation in government contexts outside the Global North. While much of the existing literature on algorithmic governance draws on cases from the United States and Western Europe, criminal justice institutions in Latin America face distinct challenges related to institutional capacity, digital infrastructure, and legal frameworks. By documenting the design choices and implementation experience in a Mexican prosecutor’s office, we offer insights that may inform similar efforts in other resource-constrained settings.
The prosecutorial backlog in Mexico’s criminal justice system poses a major institutional challenge. Across the country, prosecutors’ offices are responsible for overseeing every stage of criminal proceedings—opening investigations, collecting evidence, promoting alternative solutions, and pursuing criminal action in court. However, growing caseloads and limited resources have led to large volumes of unresolved cases. Between 2018 and 2022, the number of open files nationwide increased by nearly 60%, with over 2.9 million active investigations reported by the end of 2022 México Evalúa (2023); Instituto Nacional de Estadística y Geografía (INEGI) (2023).
In this project, we partnered with the Fiscalía General de Justicia del Estado de Zacatecas to design and evaluate a machine learning (ML) system to support weekly case prioritization. Our aim is not to automate decisions or restructure workflows, but to assist prosecutors by providing a ranked list of cases that, based on historical patterns of case progression, are more likely to be resolved in the near term.
The lists generated by the models serve as a support tool, helping the prosecutor’s office workers keep potentially actionable cases visible and accessible, without changing institutional procedures or increasing staff workload. The goal is to enable more informed prioritization while staying within current operational capacity.
In addition to identifying cases likely to be resolved in the short term, the system also helps manage the backlog by highlighting cases at the tail of the distribution. Using the same prediction framework, we generate weekly lists of approximately 1000 investigations with the lowest estimated probability of concluding within the next six months. Historical patterns indicate that many of these cases have already surpassed statutory deadlines for prescription or show signs of prolonged inactivity.
This paper presents the design and historical validation of this ML-based prioritization system. We focus on the Módulo de Atención Temprana (MAT), which handles intake and early-stage filtering. Using data from the Plataforma de Integración de Expedientes (PIE)—Zacatecas’s official case management database—we train a set of classification models to predict short-term resolution probabilities and evaluate their performance using precision-focused metrics aligned with institutional constraints.
We show that ensemble classifiers, especially Random Forests, can reliably identify cases likely to finalize within a 6-month horizon, offering a low-overhead decision-support tool for backlog management. The results from retrospective validations suggest that such models can be feasibly integrated into prosecutorial workflows to improve case triaging without disrupting operations.
2 Institutional Context: The Zacatecas Prosecutor’s Office
The Fiscalía General de Justicia del Estado de Zacatecas (FGJEZ) is responsible for overseeing the criminal investigation process across the state. In 2021, it began implementing a strategic criminal prosecution policy aimed at reducing procedural delays and improving efficiency. As part of this effort, the office has adopted a structured case management approach that distributes investigations across specialized units and, more recently, has begun integrating predictive tools to support internal prioritization. At the core of this structure is the “Modelo de Tres Pisos” (Three-Tier Model), which assigns cases to different prosecutorial units based on procedural stage and case complexity:
  • Tier 1: Módulo de Atención Temprana (MAT) — Responsible for initial complaint intake, early-stage investigation, and rapid resolution, either through alternative mechanisms or administrative closure.
  • Tier 2: Unidad de Tramitación Masiva de Casos (UTMC) — Focuses on cases with identified suspects, emphasizing resolution via conditional suspensions or abbreviated procedures.
  • Tier 3: Unidad de Investigación Especializada (UIE) — Handles complex or sensitive cases requiring full investigation and trial preparation.
Two additional support units work alongside this tiered structure: the Órgano Especializado en Mecanismos Alternativos (OEMASC), which facilitates out-of-court settlements, and the Unidad de Imputado Desconocido (UID), which handles cases without identified suspects. The main analysis in this paper focuses on the MAT unit; Appendix C reports results for applying the same methodology to UID.
As of August 2024, the PIE case management system contained over 234,000 investigation files. Roughly 55,000 remained open from previous years. The system registers an average of 1,831 new cases per month, while only 746 are formally closed, and 217 are resolved through alternative mechanisms. These figures, drawn from administrative data covering 2014 to 2024, highlight the backlog challenges faced by entry-point units like MAT, where high inflow and limited throughput generate persistent caseload pressure. To support internal case review, this paper presents a predictive prioritization system designed for the MAT unit. Each week, the model produces a ranked list of 300 open cases that, based on historical patterns, resemble files that were previously resolved within a six-month window. We share these lists with unit staff to inform daily decision-making, without modifying formal assignments, workloads, or procedural discretion. The objective is not to change how prosecutors work, but to make potentially resolvable cases more visible—supporting more effective triage within existing capacity and institutional constraints.
3 Data
The primary dataset used in this analysis comes from the Plataforma de Integración de Expedientes (PIE), the case management system maintained by the Zacatecas State Prosecutor’s Office. Since its implementation in January 2014, PIE has recorded every investigation initiated across all jurisdictions in the state, serving as the official digital ledger of prosecutorial activity. Each case file within PIE contains structured metadata, entered by prosecutors, investigators, and administrative staff through a secure web interface at key procedural milestones.
PIE captures a broad range of information for each investigation. The information includes fundamental details such as the date and location of the reported offense, the individuals involved, applicable legal classifications, and a brief narrative summary of the alleged events. Additionally, the system logs a chronological sequence of procedural events, including arrests, evidence submissions, protective measures, unit transfers (with reasons and timestamps), court referrals, and final case outcomes (e.g., closures or jurisdictional transfers). While the complete paper dossiers remain offline, PIE preserves the structured fields necessary to reconstruct the full procedural history of each case.
To support data integrity, the PIE IT team performs quarterly audits comparing digital entries to physical case files, and unit supervisors conduct weekly spot checks to monitor for delayed or missing updates. Despite these governance practices, two important limitations remain. First, the current PIE system does not digitize or allow uploading of narrative documents and evidentiary attachments—which could provide deeper insights into case complexity. Second, staff sometimes enter updates in batches, leading to timestamp noise of up to one week in some event records. To address this, we consistently rely on the update timestamps recorded in PIE rather than auxiliary time indicators to preserve consistency in event sequencing.
4 Features
Using data from the Plataforma de Integración de Expedientes (PIE), we construct a set of features designed to capture how the Prosecutor’s Office interacts with each case over time. Our unit of prediction is the individual case record, specifically those that remain open in the Módulo de Atención Temprana (MAT) at the time of analysis. The features span both static characteristics of the case and dynamic indicators of prosecutorial activity, reflecting not only the nature of the crime but also the administrative and legal progression of the case within the office. This strategy aligns with recent work emphasizing event-driven features and longitudinal context in high-stakes domains Chouldechova et al. (2018); Kleinberg et al. (2018).
We organize the features into five main groups:
  1. 1.
    Case-Level Characteristics: These features describe the fundamental attributes of each case:
    • Opening details: date the case was opened, principal crime category (categorical), and whether the case began with an arrested suspect (typically when the police initiate the case with a detained individual, often caught in flagrante).
    • We define a procedural event as any logged activity reflecting meaningful progress or change in a case’s status. PIE tracks the following core event types: case initialization, progress updates in the investigation, transfers between units or jurisdictions, issuance or execution of search warrants, updates to suspect information, updates to the involved party information, and case closure or resolution.
    • Event dynamics: Using these event types, we construct cumulative counts of total events per case, as well as temporal aggregates over the past 3 months, 6 months, 1 year, and 2 years. We also calculate average event rates per period.
    • Temporal gaps: number of days between the date of the crime and the date it was reported, days since the last recorded event, days the case has remained with the same lawyer, and the total duration since the case was opened.
  2. 2.
    Investigation Milestones: This group captures key investigative and judicial actions that signal progression or resolution pathways within a case. These include binary indicators and cumulative counts of events such as judicialization, imposition of preventive detention, requests for judicial authorization, issuance of arrest warrants, referrals to conciliation, linkage to trial (vinculación a proceso), and the use of alternative justice mechanisms. For each milestone type, we compute both historical totals and time-bounded aggregates over recent periods, enabling temporal sensitivity in predictive patterns.
  3. 3.
    Lawyer Activity and Performance: To reflect the prosecutorial dimension of each case, we engineer features that measure both individual lawyer activity and their broader engagement patterns. We record how many different prosecutors have handled a case, and for each assigned lawyer, we calculate their active caseload at the time of the case’s most recent event. Historical activity metrics include total opened, closed, and transferred cases, along with productivity ratios such as the share of open cases that have been successfully closed. We compute these over the last month, 3 months, 6 months, and full case history. We also include relative performance rankings among prosecutors, enabling the model to adjust for workload heterogeneity across staff.
  4. 4.
    Unit-Level Indicators: In parallel with lawyer-specific features, we include metrics that capture performance and workload at the unit level. These include total open and closed cases handled by each unit, recent closure rates, and activity-based rankings relative to other units. These indicators provide a contextual backdrop for understanding how institutional bottlenecks or momentum may shape case trajectories.
  5. 5.
    Crime-Type Aggregates: Finally, we include features summarizing historical trends within each crime type. For each primary offense category, we compute historical and recent counts of open, closed, and total cases, average event volumes, and closure ratios. Rank-ordering these categories by resolution activity further allows the model to recognize patterns in how different types of crime are typically processed over time.
This feature set reflects principles from interpretable machine learning design, prioritizing temporally-grounded and institutionally meaningful variables Lakkaraju et al. (2016); Molnar (2020). By focusing on longitudinal event structures rather than static case attributes, the model captures not just what kind of case is being handled, but how it is evolving—by whom, for how long, and with what apparent momentum.
5 Methodology
We frame the problem as a binary classification task: given an open case in the target prosecutorial unit (MAT), predict whether the case will be finalized within the next six months. The aim is to support the unit in prioritizing cases that are most likely to leave its jurisdiction—either because no further prosecutorial action is feasible or because the case is transferred to another unit for continued processing. Model training uses historical case data from across the entire prosecutor’s office, allowing the classifier to learn patterns of case progression observed throughout the institution. Predictions and prioritization are restricted exclusively to cases currently handled by the MAT unit.
To simulate real-time deployment, we use a rolling temporal cross-validation strategy Roberts et al. (2017). At each prediction point, we train the model only on data available up to that moment, ensuring no information leakage into the future. This approach has been recommended for high-stakes public sector applications where real-time decision support must mirror deployment conditions.
Model training uses case histories from 2014 to 2024, with predictions and evaluation focused on the 2023–2024 period. We evaluate models using Precision@300 and Recall@300, which are operationally relevant metrics aligned with institutional constraints, following similar practices in ML-for-policy research Lakkaraju et al. (2016); Jung et al. (2017).
We compare several classification models, including Logistic Regression, Decision Trees, Random Forests, and Extra Trees. In addition to these learning algorithms, we use two points of comparison to contextualize model performance: (1) a random selection of 300 cases, and (2) a base rate ranking derived from historical closure probabilities—i.e., the empirical likelihood that similar cases have been finalized in the past. While the first serves as a naive baseline, the second provides a data-informed benchmark against which to assess whether the machine learning models learn patterns beyond historical averages (see Appendix A).
Overall, this methodology aligns predictive modeling with institutional practice—supporting real-time decision-making, avoiding temporal leakage, and evaluating performance through actionable, interpretable metrics.
A key design principle underlying this system is that algorithmic tools in government should serve multiple governance functions when possible. Rather than building separate systems for productivity enhancement and legal oversight, we intentionally designed a unified framework that addresses both needs from a single prediction model. The same probability estimates that identify cases likely to be resolved in the near term also reveal, at the opposite tail of the distribution, cases that appear stalled and may warrant administrative review. This dual-purpose architecture reflects an emerging principle in public sector AI: that well-designed algorithmic systems can simultaneously support operational efficiency and institutional accountability (Peeters and Widlak, 2018; Veale et al., 2018).
In addition to the primary task of predicting short-term case resolution, we use the same prediction outputs to identify cases at high risk of statutory prescription. The need to facilitate administrative and legal oversight of long-inactive cases that are unlikely to make progress under current circumstances drives this extension. However, it may still require a formal review to determine whether these cases should be closed due to statutory prescription. By using the same model, we can maintain a consistent ranking of cases based on their anticipated future outcomes. This approach allows us to prioritize near-term actions while also assessing long-term dormancy, all derived from a single, coherent representation of case dynamics. Exact legal computation of prescription is not possible with the available administrative data, as the database lacks the legally necessary qualifiers to determine the precise offense subtype, the legally relevant starting point of the limitation period, and formal procedural acts that interrupt the limitation period. We therefore construct an operational approximation grounded explicitly in the prescriptive rules of the Código Penal para el Estado de Zacatecas (CPZac, arts. 93–112).
Across the dataset, investigations fall into 183 unique crime categories. However, under CPZac, each category typically contains multiple legally defined subtypes with distinct penalty ranges and corresponding statutory prescription periods. For example, theft may vary by the monetary amount, the use of violence, and aggravating factors, each of which implies a different prison sentence and therefore a different limitation period. To address this legal heterogeneity, we compute statutory prescription periods directly from the penalty structure of CPZac using the legally defined rules:
  • If the offense is punishable only by a fine, the prescription period is set to 1 year (art. 97 CPZac).
  • If the offense is punishable only by disqualification, suspension, or deprivation of rights, the prescription period is set to 2 years (art. 99 CPZac).
  • If the offense carries a prison sentence (either exclusively or), the prescription period equals the arithmetic mean of the minimum and maximum prison term (término medio aritmético, TMA), with a statutory lower bound of 3 years (art. 98 CPZac).
  • Certain sexual offenses involving minors or legally incapable victims are legally imprescriptible (art. 96, second paragraph CPZac) and are therefore excluded from prescription screening.
Because each crime category $c$ may map to multiple legal subtypes with different statutory penalties, we construct three alternative statutory prescription thresholds for each category. Let $\mathcal{S}(c)$ denote the set of all legally defined subtypes associated with crime $c$, and let $T(s)$ denote the statutory prescription period implied by subtype $s\in\mathcal{S}(c)$ under the rules above. We then define:
$T_{\min}(c)=\min_{s\in\mathcal{S}(c)}T(s),\quad T_{\text{mean}}(c)=\frac{1}{|\mathcal{S}(c)|}\sum_{s\in\mathcal{S}(c)}T(s),\quad T_{\max}(c)=\max_{s\in\mathcal{S}(c)}T(s).$
The maximum threshold corresponds to assuming the legally most severe subtype within the category (longest prison sentence), the minimum threshold corresponds to the least penalized subtype, and the mean threshold represents an intermediate statutory approximation across all lawful variants.
For each case $i$ associated with crime $c_{i}$, with opening date $t_{i}$ and evaluation date $t$, we flag a case as potentially prescribed under threshold $k\in\{\min,\text{mean},\max\}$ if:
$t-t_{i}\geq T_{k}(c_{i}).$
At each weekly prediction point, we flag cases as potentially prescribed when they simultaneously fall within the bottom tail of the predicted resolution probability distribution (lowest-ranked cases) and exceed these statutory time thresholds. This procedure does not aim to establish legal prescription with certainty, as it does not reconstruct interruptive acts or tolling events (arts. 102–104 CPZac), but rather to identify files with a high likelihood of having surpassed legally meaningful time horizons under plausible statutory assumptions. This approach narrows administrative review to cases that are both unlikely to progress and plausibly affected by statutory time limits, rather than determining prescription outcomes directly.
In the main Results section, we report estimates based on the mean statutory threshold $T_{\text{mean}}$ as our primary specification. This choice provides a balanced approximation that avoids the overly conservative assumptions of the maximum threshold and the potentially permissive assumptions of the minimum threshold. We report results based on the full range of thresholds in the Appendix as robustness checks.
6 Results
Unless otherwise noted, all results in this section refer to cases handled by the MAT unit. Appendix C summarizes experiments for the Unidad de Imputado Desconocido (UID).
Figure 1 presents model performance over time, showing that ensemble methods—particularly Random Forests—consistently outperform simpler baselines. This aligns with findings from other ML applications in legal and public service domains, where ensemble classifiers have proven effective at detecting latent patterns in administrative data Berk et al. (2018); Wang et al. (2020).
Figure 1: Precision@300 for best model per algorithm across evaluation periods.
As shown, both the Random Forest and Extra Trees classifiers consistently outperform other algorithms over time. While their performance is closely aligned, the Random Forest slightly edges out the Extra Trees on average, making it the top-performing algorithm overall. Simpler models like the Decision Tree and Scaled Logistic Regression also perform reasonably well, but with greater variability across time. The Dummy Classifier, which randomly selects 300 cases, provides a reference point for how much signal the predictive models capture beyond chance. In contrast, the empirical baseline—constructed by ranking cases by historical closure rates, grouped by principal crime, and estimated from the full prosecutor’s office data—offers a data-driven benchmark that reflects past institutional patterns rather than random selection.
Beyond retrospective evaluation, we evaluated the three best-performing Random Forest models on newly available data extending through January 2025 using the same rolling temporal framework. Figure 2 reports the Precision@300 of these models over the extended horizon. Performance remains stable relative to prior evaluation periods, indicating that the models generalize well under real-time conditions and sustain their ability to surface high-likelihood cases over time.
Figure 2: Out-of-time Precision@300 of the best-performing Random Forest models evaluated on data through January 2025.
Table 1 summarizes the average Precision@300 for the best model of each algorithm over the full evaluation window.
Table 1: Average Precision@300 for the best model per algorithm across evaluation periods
Model Type Average Precision@300
Random Forest Classifier 0.735
Extra Trees Classifier 0.691
Scaled Logistic Regression 0.618
Decision Tree Classifier 0.573
Dummy Classifier 0.258
Empirical Baseline 0.253
Overall, these results suggest that ensemble-based classifiers—particularly Random Forests—are most effective at identifying high-likelihood case completions under real-time constraints. The consistent gap between these models and both the simpler classifiers and the random baseline highlights the utility of learning from detailed prosecutorial event data.
6.1 Feature Importance
To understand what drives the model’s predictions, we analyzed feature importances from the top-performing model group: the Random Forest classifiers. Figure 3 shows the aggregated importance scores for different feature groups, averaged across evaluation periods. We compute these group-level importances by summing the normalized importance of all individual features within each group, providing a higher-level view of where predictive power resides.
Figure 3: Aggregated feature importance over time by feature group for Random Forest models.
As shown, the most influential group of features corresponds to procedural movements (moves), which include event-level information such as the number and timing of case status changes, unit transfers, or investigation milestones. This category consistently dominates across all periods, suggesting that case activity patterns offer strong predictive signals of whether a case is nearing resolution. Other important groups include days_since_open (how long the case has been active), days_since_last_event (recency of prosecutorial action), and estados_investigacion (status transitions in the investigation).
To drill down into the specific variables contributing most to prediction, Figure 4 presents the top 20 individual features ranked by importance. The strongest single predictor is the number of "initiated" events in the past year, followed closely by the same feature measured over six-month and two-year windows. In this context, "initiated" refers not to a case being reopened after closure, but rather to a formal re-initiation event caused by a transfer between units, which triggers a new entry in the PIE system. These events reflect administrative handoffs, signaling that a case has been reassigned and is now actively progressing under a different team. The frequency of these transitions may indicate cases that are moving through the system toward resolution.
Other high-importance features include the minimum number of days since case opening, as well as rolling averages of specific event types such as avance_investigacion (investigation progress) and cambio_unidad (unit transfers). These temporal dynamics collectively suggest that recent and sustained prosecutorial activity—especially related to reassignment and procedural advancement—are key indicators of eventual case completion.
Figure 4: Top 20 most important features in the best-performing Random Forest models.
The model primarily depends on dynamic signals that capture how cases evolve—particularly the frequency and timing of administrative or investigative actions. In contrast, static attributes such as the presence of an arrested suspect at intake or generalized crime type statistics contribute less to prediction. These results underscore the importance of the longitudinal event structure captured in PIE and demonstrate the value of modeling prosecutorial workflows over time, rather than relying solely on initial case characteristics.
6.2 Oversight Function: Identifying Potentially Prescribed Cases
In addition to prioritizing cases with a high likelihood of near-term resolution, we analyze the opposite tail of the prediction distribution by examining the 1000 cases with the lowest predicted probability of concluding within six months. We evaluate these cases for potential statutory prescription using the three operational definitions described in the Methodology. In this section, we report results based on the mean statutory prescription threshold $T_{\text{mean}}$, while we present the minimum and maximum thresholds in Appendix B as robustness checks.
Table 2 reports the mean number and corresponding share of potentially prescribed cases among the bottom 1000 for the best-performing model of each algorithm under the mean statutory threshold. Under this specification, the Random Forest model flags, on average, 318.6 cases as potentially prescribed, corresponding to 31.9% of the low-priority tail. The Extra Trees and Scaled Logistic Regression models identify approximately 212–213 cases (about 21.3%), while the Dummy Classifier flags 230.3 cases (23.0%).
Table 2: Potentially prescribed cases among the 1,000 lowest-ranked cases (mean statutory threshold)
Model Mean Count Share (%)
Random Forest 318.63 31.9
Extra Trees 215 21.5
Scaled Logistic Regression 213.43 21.3
Dummy Classifier 50.95 5.1
Figure 5 shows the temporal evolution of the mean number of potentially prescribed cases among the bottom-ranked files under the mean statutory threshold for all model families. Across all model types, the number of flagged cases exhibits a persistent upward trend over time. This pattern indicates that prescription risk in the low-priority tail is not episodic but instead accumulates steadily as cases remain inactive.
Importantly, the trend is qualitatively similar across the Random Forest, Extra Trees, and Scaled Logistic Regression models, suggesting that the observed increase reflects structural backlog dynamics rather than model-specific trends. The Dummy Classifier also displays a rising trajectory, reinforcing the interpretation that prescription accumulation is driven primarily by institutional time dynamics rather than predictive ranking alone.
Figure 5: Mean number of potentially prescribed cases among the bottom 1000 cases under the mean statutory threshold, by model family.
To assess robustness within the top-performing Random Forest family, Figure 6 reports the same prescription dynamics for the three best Random Forests. Across these configurations, the prescription signal is highly stable: the mean number of potentially prescribed cases in the bottom 1000 remains tightly concentrated around 316-319 cases over time. This stability indicates that the prescription identification is not sensitive to hyperparameter variation within the Random Forest class.
Figure 6: Mean number of potentially prescribed cases among the bottom 1,000 using the mean statutory threshold for the three best Random Forest models.
Taken together, these results indicate that roughly one-third of the lowest-priority cases identified by the best-performing models are likely to have surpassed legally meaningful statutory time limits under the mean prescription rule. This additional list of the lowest-ranked cases complements the institutional function of the prioritization system. While the top of the ranking supports short-term prosecutorial productivity, the bottom supports a systematic, data-driven mechanism for identifying files that warrant administrative review, formal closure, or corrective action.
7 Implementation and Future Steps
While the selected model shows strong performance on historical data, we need to validate its effectiveness in a real-time operational setting before integrating it into prosecutorial workflows. To understand the real-world impact of the model and the prioritization system it supports, we plan to conduct a randomized controlled trial (RCT) in the Módulo de Atención Temprana (MAT). This evaluation aims to determine whether prioritizing cases based on the model’s output leads to measurable improvements in prosecutorial throughput under routine operating conditions Kleinberg et al. (2018).
The selected model—a Random Forest classifier—achieved an average Precision@300 of 0.74 across historical validation periods. While this indicates strong discriminatory performance, retrospective accuracy alone cannot establish real-time utility. We structure the RCT to causally evaluate whether model-informed prioritization increases the number of cases finalized or reduces time to resolution relative to the status quo Chouldechova et al. (2018).
In the proposed experimental design, we will run the model weekly to score all open cases in the MAT unit. We will identify the 600 cases with the highest predicted probability of being finalized within the next six months. From this ranked set, we will randomly assign 300 cases to the treatment group and mark them for prioritized review, while the remaining 300 will form the control group and proceed through the standard workflow. Prosecutors will not be informed of group assignments, and we will track case outcomes using unique administrative identifiers.
To avoid repeated exposure and dependence across weekly cohorts, we enroll each case at most once: if a case selected in a prior week remains open and reappears in the top-ranked set, we exclude it from re-randomization and replace it with the next-highest-ranked eligible case, ensuring each weekly cohort contains 600 unique cases.
We will conduct the experiment over six consecutive weeks, yielding approximately 1800 treatment cases and 1800 control cases. We will follow all cases for six months and compare outcomes along operationally relevant dimensions such as resolution rates and time to completion Chouldechova et al. (2018). This design tests the effect of prioritized visibility itself, rather than any procedural or staffing change.
In parallel with the productivity-focused RCT, we plan a second evaluation track focused on the administrative and legal implications of the system’s prescription-screening component. Each week, the model identifies a subset of low-priority cases that exceed the mean statutory prescription threshold. Although this signal does not constitute a formal legal determination, it provides a data-driven list of files warranting administrative review.
Because there is no single outcome metric analogous to resolution for prescription, evaluation will rely on process-based indicators. Specifically, we will track the actions taken following the recommendation (e.g., formal verification of prescription, decision not to prosecute archival, reopening due to new information, or continued investigation), the time elapsed between the recommendation and the administrative decision, and the share of reviewed cases in which the prescription is formally confirmed. This will allow us to assess whether the system facilitates earlier identification and handling of legally dormant files, even without a single summary performance statistic.
If both the prioritization and prescription components demonstrate operational value, we may incorporate the system into the MAT unit’s routine workflow with two complementary outputs: a high-priority list for near-term prosecutorial action and a low-priority list for administrative and legal review.
Finally, the model architecture and feature design developed in this study are tightly coupled to Zacatecas’s institutional processes, digital infrastructure, and historical data availability. As a result, the model is not directly transferable to other jurisdictions without substantial adaptation. Differences in case management systems, event taxonomies, legal procedures, and digitization practices would require redesigning both the features and the validation strategy. Nevertheless, the broader methodological framework—framing triage and legal dormancy as prediction problems and validating them through policy-aligned experiments—may be portable to other prosecutorial settings with compatible administrative data.
8 Discussion
This case study offers several insights for the design and governance of AI systems in public sector contexts. First, the dual-purpose architecture demonstrates that algorithmic tools can serve both operational efficiency and institutional oversight objectives simultaneously. By designing a single prediction framework that supports near-term prioritization at one tail of the distribution and statutory review at the other, we avoided the need for separate systems while ensuring that both governance functions receive attention. This approach aligns with calls for AI systems that embed accountability mechanisms from the design stage rather than treating oversight as an afterthought (Veale et al., 2018; Peeters and Widlak, 2018).
Second, the implementation experience highlights the importance of institutional fit. The system was designed around existing workflows rather than requiring organizational restructuring. Weekly lists are delivered through channels already familiar to prosecutors, and the ranked format aligns with how staff already conceptualize case prioritization. This “low-overhead” approach, which aims to enhance rather than replace professional judgment, appears essential for adoption in resource-constrained settings where staff have limited capacity to learn new systems or adapt to changed procedures.
Third, the case illustrates how administrative data infrastructures shape the possibilities for algorithmic governance. The Plataforma de Integración de Expedientes (PIE) provides the digital foundation that makes this kind of analysis possible, but its structure also constrains what can be predicted. The system captures procedural events but not the substantive content of case files; it records timestamps but with some noise; it classifies crimes at a level of aggregation that introduces uncertainty into prescription calculations. These limitations are not unique to Zacatecas—they reflect the broader reality that algorithmic systems in government must work with administrative data designed for other purposes (Janssen and Kuk, 2016).
Finally, this work contributes to the emerging literature on AI in government contexts outside the Global North. The challenges faced by the Zacatecas prosecutor’s office—high caseloads, limited staff, persistent backlogs—are common across Latin American criminal justice systems and in many other developing country contexts. By documenting both the potential and the limitations of algorithmic decision support in this setting, we hope to inform future efforts to design AI systems that are appropriate for the institutional realities of diverse governance contexts.
9 Ethical Considerations and Limitations
We designed the prioritization model presented in this paper strictly as a decision-support tool, not as a substitute for prosecutorial judgment. It does not alter decision-making authority, interfere with case assignment, or impose mandatory action thresholds. Instead, it surfaces cases that resemble past resolutions under existing institutional behavior to assist internal triage within capacity constraints. Prosecutors retain full discretion over whether and how to act on model outputs.
However, any model trained on historical administrative data may reflect patterns of past bias or structural inequality. For example, the prioritization logic may implicitly favor case types, jurisdictions, or procedural pathways that have historically moved through the system more quickly. This could unintentionally reinforce disparities if not carefully monitored Barabas et al. (2018); Kleinberg et al. (2018).
The prescription-screening component introduces additional ethical and legal sensitivities. The system does not determine legal prescription, nor does it reconstruct interruptive procedural acts or tolling events required for a formal legal declaration. Instead, it produces a risk-based shortlist of cases that appear to exceed statutory time horizons under operational legal approximations. There is a nontrivial risk that such signals could be misinterpreted as definitive legal conclusions if safeguards are not maintained. For this reason, we frame all prescription-related outputs as recommendations for administrative review, not as binding legal determinations.
To mitigate broader algorithmic risks, we implemented the system in a non-intrusive manner: prosecutors retain complete discretion, and no decisions are made automatically. Moreover, the ongoing randomized controlled trial will allow us to test whether model-guided prioritization has differential impacts across subgroups.
We conduct an initial fairness assessment to evaluate whether the model systematically prioritizes certain types of crime or unintentionally deprioritizes cases across municipalities or population groups. The analysis examines subgroup precision, recall, and exposure rates, together with potential disparities in prioritization and prescription flagging. In preliminary fairness diagnostics, we examine subgroup performance using historical prediction runs. At the municipal level, exposure and Precision@300 vary across jurisdictions: the reference municipality (Zacatecas), which also accounts for the largest share of cases, exhibits comparatively stable parity, while smaller municipalities show greater variability. By contrast, crime-type analyses look more balanced overall, with most categories tracking the reference offense (simple theft) closely, and clearer deviations appearing mainly in property-damage cases. These patterns likely indicate that differences in caseload structure and institutional processing are responsible, but they also underscore the need to monitor fairness rather than assume it empirically.
We will continue expanding these diagnostics in future iterations, incorporating additional fairness metrics and stability checks as more deployment data becomes available.
Finally, because we approximate statutory prescription using generic crime categories rather than legally exact offense subtypes, all prescription-related estimates must be interpreted as upper- and lower-bound risk indicators rather than precise legal conclusions. This limitation reinforces the need for human legal verification in all downstream uses of the system.
Data Availability
The data used in this study contains sensitive criminal justice information and cannot be publicly shared due to legal and privacy restrictions under Mexican law. The administrative records from the Plataforma de Integración de Expedientes (PIE) include personal identifiers, crime classifications, and case disposition information that are protected under federal and state data protection regulations. Aggregated results, model specifications, and additional methodological details are available from the corresponding author upon reasonable request.
Declaration of Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. The collaboration between Tecnológico de Monterrey and the Fiscalía General de Justicia del Estado de Zacatecas was conducted as part of ongoing institutional cooperation on public sector innovation.
Author Contributions (CRediT)
Fernanda Sobrino: Methodology, Software, Formal analysis, Writing – original draft, Visualization. Adolfo De Unánue: Conceptualization, Methodology, Supervision, Writing – review & editing. Edgar Hernández: Data curation, Software. Patricia Villa: Investigation. Elena Villalobos: Visualization, Data curation. David Aké: Project administration. Stephany Cisneros: Investigation. Cristian Paul Camacho Osnay: Conceptualization, Resources, Project administration. Armando García Neri: Conceptualization, Supervision, Resources. Israel Hernández: Data curation, Resources.
Appendix A Construction of the Historical Base-Rate Baseline
This appendix describes how we construct an empirical baseline to contextualize the performance of the machine learning models. The baseline uses a non-machine-learning prioritization rule and relies only on historical case outcomes. We evaluate it under the same rolling temporal framework and outcome definitions as the predictive models.
At each evaluation week $t$, we identify all open and active cases in the MAT unit. To estimate historical closure rates, we use all case outcomes recorded across the prosecutor’s office prior to week $t$. This broader historical sample allows us to estimate stable baseline closure patterns while preserving the temporal separation between training and evaluation.
We group cases by their principal crime category as recorded in the administrative system. For each crime group $k$, we estimate the probability that a case in that group will be finalized within 6 months. We compute these probabilities using all historically observed cases in the group with observation dates strictly earlier than $t$. Each historical label indicates whether the case was finalized within six months of its observation date.
When a crime group contains limited historical data, we smooth its estimated closure probability toward the overall historical closure rate using an additive shrinkage approach. This procedure balances group-specific information with aggregate historical trends and reduces variance caused by small sample sizes.
At week $t$, we assign each open MAT case the estimated six-month closure probability corresponding to its crime group. We then rank cases in descending order of this score, select the top 300 cases, and compute Precision@300 using the same evaluation windows and outcome definitions as the machine learning models.
This baseline uses only information available prior to the prediction date and incorporates no case-level features beyond crime category. It therefore reflects historical institutional patterns rather than learned predictive relationships, providing a transparent and interpretable point of comparison for the machine learning results.
Appendix B Sensitivity to Alternative Statutory Prescription Thresholds
This appendix reports results using the minimum and maximum statutory prescription thresholds defined in Section 5, in addition to the mean threshold used in the main text. Here we summarize how the number of potentially prescribed cases in the bottom 1000 changes under the more permissive ($T_{\min}$) and more conservative ($T_{\max}$) thresholds.
B.1 Best Model per Algorithm
Table 3 reports the mean number of potentially prescribed cases among the 1000 lowest-ranked files for the best-performing model of each algorithm, under each of the three thresholds.
Table 3: Mean number of potentially prescribed cases in the bottom 1,000 under alternative statutory thresholds
Model $T_{\max}$ $T_{\text{mean}}$ $T_{\min}$
Random Forest 109.52 318.63 651.78
Extra Trees 118.53 212.48 643.45
Scaled Logistic Regression 95.80 212.70 663.85
Dummy Classifier 31.28 50.95 138.78
As expected, the conservative threshold $T_{\max}$ yields the smallest number of flagged cases (roughly 9–12% of the bottom 1,000), while the permissive threshold $T_{\min}$ yields the largest (around 64–75%). The mean threshold $T_{\text{mean}}$ lies between these extremes, flagging roughly 21–32% of the low-priority tail depending on the model. The Random Forest model remains the most effective at concentrating potentially prescribed files in the bottom rankings across all three thresholds.
B.2 Best Random Forest Model Groups
Table 4 presents the same sensitivity analysis for the three best Random Forest model groups (IDs 32, 37, and 38).
Table 4: Mean number of potentially prescribed cases in the bottom 1,000 for the three best Random Forest models
RF Model Group $T_{\max}$ $T_{\text{mean}}$ $T_{\min}$
RF 1 106.49 316.07 654.75
RF 2 109.52 318.63 651.78
RF 3 108.92 318.83 659.27
Across these top Random Forest configurations, the number of potentially prescribed cases is very similar under each threshold, with variation of only a few cases around the mean-rule value of approximately 318 files. This suggests that the prescription screening results are robust to reasonable changes in Random Forest hyperparameters as well as to the choice of statutory threshold within the legally admissible range.
Appendix C Results for the Unidad de Imputado Desconocido (UID)
In addition to the main analysis for the Módulo de Atención Temprana (MAT), we conducted experiments applying the same modeling framework to the Unidad de Imputado Desconocido (UID), which handles investigations without identified suspects. UID cases differ substantively from MAT cases: they typically enter the system with less information about potential perpetrators, progress more slowly, and often remain open for extended periods with limited procedural activity.
For UID, we replicate the MAT setup as closely as possible. We restrict the prediction sample to cases that are open in UID at each evaluation week, define the outcome as whether the case finalizes within six months, and use the same feature construction, rolling temporal evaluation, and model class (Decision Tree, Random Forest, Extra Trees, Scaled Logistic Regression), together with the empirical and random baselines.
Figure 7 reports Precision@300 over time for the best-performing model of each algorithm in UID. In contrast to the MAT results, the models achieve substantially lower precision throughout the evaluation window, and the gap between the machine learning models and the empirical baseline is modest. This pattern is consistent with the underlying institutional dynamics of UID: very few cases without identified suspects resolve within a short horizon, and the available administrative signals provide limited information about which of these cases will close next.
Figure 7: Precision@300 for the best model per algorithm in UID across evaluation periods.
Overall, these results suggest that the short-term completion prediction task is harder in UID than in MAT, and that the current feature set and label definition provide limited leverage for prioritizing “next-to-finish” cases when suspects remain unidentified.
We also apply the prescription-screening procedure described in Section 5 to the UID setting. As in MAT, we examine the 1000 cases with the lowest predicted probability of being concluded within six months and evaluate them under the mean statutory prescription threshold, $T_{\text{mean}}$.
Figure 8 shows the temporal evolution of the mean number of potentially prescribed cases among these bottom-ranked UID files. Across all model families, a large share of low-priority UID cases exceeds the mean statutory threshold, and the number of flagged cases remains high and relatively stable over time. This pattern reflects the fact that UID predominantly hosts long-running investigations with limited recent activity, which are structurally unlikely to progress but may nonetheless remain open in the administrative system.
Figure 8: Mean number of potentially prescribed cases among the bottom 1,000 UID cases under the mean statutory threshold, by model family.
To assess the sensitivity of these results to the choice of statutory threshold, Table 5 reports the mean number of potentially prescribed cases among the bottom 1000 UID files under the mean, maximum, and minimum prescription rules. For each model, $T_{\text{mean}}$ flags roughly half to over ninety percent of the lowest-ranked UID cases as potentially prescribed, while the more conservative $T_{\max}$ flags only around 4–9% of cases and the more permissive $T_{\min}$ flags roughly 72–97%.
Table 5: Mean number of potentially prescribed UID cases in the bottom 1000 under alternative statutory thresholds
Model $T_{\text{mean}}$ $T_{\max}$ $T_{\min}$
Decision Tree Classifier 734.13 77.75 863.80
Random Forest Classifier 780.28 71.63 958.63
Extra Trees Classifier 895.40 78.20 972.73
Scaled Logistic Regression 922.28 92.13 954.18
Dummy Classifier 498.53 37.30 723.40
Taken together, the UID results indicate that while predictive models add little value for identifying short-term completions in this unit, the prescription-screening component remains informative. Even when discriminative performance on the “next-to-finish” label is weak, the combination of long case durations and low predicted resolution probabilities yields a concentrated set of UID files that merit administrative and legal review for potential statutory prescription. In this sense, the prescription module provides a useful diagnostic tool for backlog governance in UID, even if prioritization for near-term resolution is less promising than in MAT.